For drivers aged 20-24 there is a 34% chance of having a car accident in a one year period (based on data from the National Safety Council).
(a) Based on this data, in a group of 14 randomly selected drivers aged 20-24, find the probability that at least 2 of them will have a car accident in the next year.

Answer by rothauserc(4718)

We use the binomial probability distribution to solve this problem
:
Probability (P) (k successes in n trials) = nCk * p^(k) * (1-p)^(n-k), where nCk = n! / (k! * (n-k)!)
:
n = 14, p = 0.34
:
P (at least 2 drivers have an accident) = 1 - P(all 14 drivers have no accidents) - P(1 driver out of 14 has an accident)
:
P (at least 2 drivers have an accident) = 1 - 0.003 - 0.021 = 0.976 approximately 0.98
: